# A Rigorous Behavior Assessment of CNNs Using a Data-Domain Sampling Regime

### v1

@@ 4.3 Results
Overview. We collected a total of 4 million predictions from 200 models in each experiment ().
Figure 5 and Table 2 show the mean errors and the summary statistics. See additional analyses in Apdx. Figure 28 and Apdx. Table 6.
Sampling method is a significant main effect.
Sampling effect.H1 was supported for height sampling but not for the ratio sampling experiment. For the ratio sampling experiment, ADV (0.013) was in the same Tukey group as COV (0.011) and IID (0.011) (Figure 5
For height sampling,
COV (0.015) and IID (0.027) had lower errors than ADV (0.042) and OOD (0.059) (Figure 5
Striking generalization failure on OOD. Sampling effect.H2 was supported. To study the CNNs’ behaviors, we analyzed case-by-case CNNs’ inference accuracy by ratio bins (Figure 6
@@ 6.3 Results
Overview. We collected human observers’ results from 1,575 trials from 63 participants, who spent an average of 10.0 seconds per trial.
Figure 10, Figure 11, and Table 4 show the mean errors and the summary statistics. Overall, our human experiment results showed slightly higher error than Cleveland and McGill’s (Figure 10
Chart type was also a significant main effect. The differences between types were observed in human experiments (Figure 10
@@ Appendix B A Comparison of Sampling Input Distributions and Data Processing Approaches
Cleveland and McGill [8] (
@@
Talbot et al. [45] (
Haehn et al. [18] (
Our replication also used the 10 integer heights (10, 12, 15, 18, 21, 26, 32, 38, 46, 56). We also removed height combinations [(10, 46), (12, 56)] and [(10, 32), (12, 38), (15, 46), (18, 56)], which will produce ratios 0.18 and 0.32 in the original Cleveland and McGill’s experiment if the heights were not rounded.

### v2

@@ 4.3 Results
Overview. We collected a total of 4 million predictions from 200 models in each experiment (). Figure 5 and Table 2 show the mean errors and the summary statistics. See additional analyses in Apdx. Figure 28 and Apdx. Table 6. Sampling method is a significant main effect. Sampling effect.H1 was supported for height sampling but not for the ratio sampling experiment. For the ratio sampling experiment, ADV (0.013) was in the same Tukey group as COV (0.011) and IID (0.011) (Figure 5 ), meaning that we did not observe a statistical difference among them. One reason for the lack of significance for ADV could be that it contained sufficient unique samples across the data domain.
For height sampling, COV (0.015) and IID (0.027) had lower errors than ADV (0.042) and OOD (0.059) (Figure 5 ). Now ADV was in a different Tukey group (Table 2). We suspected that the differences between IID, COV, and ADV conditions were caused by errors introduced by the smaller bar heights in Figure 6.
Striking generalization failure on OOD. Sampling effect.H2 was supported. To study the CNNs’ behaviors, we analyzed case-by-case CNNs’ inference accuracy by ratio bins (Figure 6 ) and by heights (Figure 6 ) computed from the two experiments correspondingly. In Figure 6 , the diagonal line means that the prediction is a perfect match to the ground truth, and the closer to this line, the smaller the inference errors. Figure 6 shows the mean absolute errors against bar heights, lower error indicates better model performance. We observe that knee-shaped curves emerge in the out-of-train regions in OOD in both ratio sampling and height sampling experiments, showing errors that increase almost linearly. Interestingly, the in-train test data points in OOD maintain low errors (0.008 for ratio sampling and 0.025 for height sampling). This result may support that CNNs’ inference accuracy depends on the coverage of input: IID did not guarantee sampling of the smallest or the largest values compared to COV and the minimum values of OOD. Our experiments in the next section validate this conjecture by stress-testing CNNs with limited training samples.
@@ 6.3 Results
Overview. We collected human observers’ results from 1,575 trials from 63 participants, who spent an average of 10.0 seconds per trial. Figure 10, Figure 11, and Table 4 show the mean errors and the summary statistics. Overall, our human experiment results showed slightly higher error than Cleveland and McGill’s (Figure 10 ), but the trend across the five chart types remained consistent with the findings of Cleveland and McGill [8] and Heer and Bostock [20] (Figure 10 ). (Apdx. Figure 32 has more details.) CNNs had lower errors than humans. Human-AI comparisons.H1 was supported. The observer was a main effect. For ratio sampling, human errors (0.056) were about four times that of CNNs’ IID-large (0.013) (Figure 10 ), and for height sampling, human errors (0.058) were about three times of CNNs’ IID-large (0.019) (Figure 11 ).
Chart type was also a significant main effect. The differences between types were observed in human experiments (Figure 10 and Figure 11 ). In addition, CNNs had smaller uncertainty (95% CI for Type 1, Baseline, ratio: 0.002, height: 0.002) compared to human observers (95% CI, ratio: 0.007, height: 0.004).
@@ Appendix B A Comparison of Sampling Input Distributions and Data Processing Approaches
Cleveland and McGill [8] ( in Figure 22). Seven ratios from a geometric sequence evenly spaced on a log scale of 10 bar heights , where (). The 45 combinations of those 10 heights led to 9 distinct ratios (0.18, 0.22, 0.26, 0.32, 0.38, 0.46, 0.56, 0.68, 0.83). 0.22 and 0.32 were not used because they were too close to 0.18 and 0.26.
@@
Talbot et al. [45] ( in Figure 22) drew bar charts on a 200px200px canvas. Their three taller bar heights were 50px, 100px, and 150px. Ratios were the same 7 ratios as Cleveland and McGill’s [8]: 0.178, 0.215 0.261, 0.316, 0.383, 0.464, 0.562, 0.682, 0.825.
Haehn et al. [18] ( in Figure 22). Used the 10 integer bar heights (10, 12, 15, 18, 21, 26, 32, 38, 46, 56). The simulation of Haehn et al.’s 12 exemplar runs in Figure 23.
Our replication also used the 10 integer heights (10, 12, 15, 18, 21, 26, 32, 38, 46, 56). We also removed height combinations [(10, 46), (12, 56)] and [(10, 32), (12, 38), (15, 46), (18, 56)], which will produce ratios 0.18 and 0.32 in the original Cleveland and McGill’s experiment if the heights were not rounded. in Figure 22 except for the excluded combinations near ratio 0.18 and 0.32 were used in our experiment.